Suppose we are looking for the point on the plane x + 2y + z = 5 closest to the point (2,3,0). Which of the following approaches DOES NOT lead to the answer? = 0 Solve the system of equationsJ 2(x - 2) - 2(5 - x - 2y) | 2(y-3) - 4(5 - x - 2y) = 0 Find the intersection point of the line r(t) = (2+ t, 3 + 2t, t) with the given plane. 2(x 2 2(y 3) = 0 = 0 Solve the system of equations 22 = 0 2( 2) 2(y 3 = > =22 Solve the system of equations 2z x 2yz= 5 Question 5 2 pts Consider the function f(x,y, z) = 5 + (x - 6)2+ (y- 2)2 + (z - 4)2 on the sphere a2 y2 + z2 = 14. Then, , and the absolute minimum is 25 (V14- 6)2 the absolute maximum is 25+(-/14 - 6)2 4-6) ( 14 - 4)2 - 2)2 + the absolute maxi mum is 5 + .and the absolute minimum is 3 (- - 6)? +(- 14 14 14 - 2)2 ( 4)2 5 1 the absolute maximum is 131, and the absolute minimum is 19. the absolute maximum is 14, and the absolute minimum is )
Consider the region D in the ry-plane bounded by the parabola y = 4- x2 and the a-axis. Consider the function f(x, y)( 1)2 2y. _ The absolute maximum value of f(x, y) for (x, y) in D is 8 The absolute minimum value of f(x, y) for (x, y) in D is 0 Question 3 3 pts Consider the function f(x,y) = 2x2 + (y - 1)2 and the disk D : x2 + y2
Answer #1

Question 5 Consider the function f(x, y, z) 5+ ( 6)2 + (y-2)2 +(z-4)2 on the sphere a2 +y+14. Then, the absolute maximum is 26 F(x y,z) 2(x-6)+ 22x =0x : 1+A 2 y2 F.(x, y.z) 2(y-2)+22y 0 4 F (x,y,z) 2(z-4)+21z 0z= 1+A 2 6 36+4+16-14(1+2 = 41+2=t2 2 4